CONJUGACY CLASSES OF THE HYPERELLIPTIC MAPPING CLASS GROUP OF GENUS 2 AND 3

KAZUSHI AHARA AND MITSUHIKO TAKASAWA

Abstract. We present tables of conjugacy classes of the hyperelliptic mapping class group of genus 2 and 3, and some theorems on the Sp representation, the Jones representation, and Meyer’s function.